How many grams of sodium fluoride form when 5.5 moles of HF reacts in this reaction?
1
Expert's answer
2013-02-26T10:27:48-0500
sodium fluoride form following the reaction:
HF+NaOH=NaF+H2O
According the equation of reaction when 5.5 moles of HF react form 5.5 moles of NaF.
We use the periodic table in order that calculate the molar mass of NaF.
M(NaF)=22.990+18.998=41.988 g/mol.
So, the mass of NaF is: m(NaF)= 41.988•5.5=230.934 g.
Answer:
The mass of NaF is 230.934 g.